OpenStudy (anonymous):

A corner stand sells peanuts for $1.00 per pound and and walnuts for $2.00 a pound. The operator of the stand wants to make 50-pound of the peanut-walnut mix. He will maintain the cost per pound of each nut and sell the mix for $1.60 a pound. How many pounds of each type of nut should he put in the mix?

OpenStudy (kirbykirby):

Let x = number of pounds of peanuts y = number of pounds of walnuts (1) Now, he is selling the mix at $1.60 per pound. That means that for 50 pounds, the mix will cost (1.6 * 50)=$80 (2) Now, you now that the total number of pounds of nuts is 50 lbs. Combining these two facts you get: sentence (1) gives you the equation \(1x+2y=80\), since a pound of peanuts is $1, and a pound of walnuts is $2. I Drop the "1" in front of \(x\) for convenience. sentence (2) gives you the equation: \(x+y=50\), the total number of pouns of both nuts is 50. So you have a system of two equations: \[\begin{cases}x+2y=80 &\\ x+y=50 &\end{cases}\] Solve for \(x\) first in both equations and equate them: \(x=80-2y\) \(x=50-y\) \(80-2y = 50 -y\) \(80-50=y\) \(30=y\) Since y=30, plug this value into one of your equations, say \(x+y=50\), \(x+30=50\) \(x=20\) \(\boxed{\text{solution: }x=20, y=30}\)